The circle (x−1)
2 +y
2 = 1,z = 0 lies inside the sphere centred at the origin, and
having radius 2√
2.
1
Expert's answer
2020-03-02T09:32:04-0500

To prove that the circle "\begin{cases}
(x-1)^2+y^2=1 \\
z=0
\end{cases}" lies inside sphere centred at the origin, and

having radius 2√2, we need to check that "x^2+y^2+z^2<(2\sqrt {2})^2" .

Proof:

"\begin{cases}
(x-1)^2+y^2=1 \\
z=0
\end{cases}, 
\quad
\begin{cases}
x^2-2x+1+y^2=1 \\
z=0
\end{cases}, \quad 
\begin{cases}
 x^2+y^2=2x \\
z=0
\end{cases}"

"x^2+y^2+z^2=2x+0=2x"

We know that "(x-1)^2+y^2=1 \ \Rightarrow (x-1)^2\leq1 \ \Rightarrow 0\leq x\leq 2"

So, we have "x^2+y^2+z^2=2x\leq 2\times 2=4<8=(2\sqrt{2})^2."

Therefore, that circle lies in the sphere.